# Characterization of birefringence inhomogeneity of KAGRA sapphire mirrors from transmitted wavefront error measurements

### v1

@@ Abstract
Future gravitational wave detectors are going to cool down their test masses to cryogenic temperatures to reduce thermal noise. Crystalline materials are considered most promising for making test masses and their coatings due to their excellent thermal and optical properties at low temperatures. However, birefringence due to local impurities and inhomogeneities in the crystal can degrade the performance of the detector. Birefringence measurement, or birefringence mapping over a two-dimensional area has become important. This paper describes a method of fast birefringence measurement for a large sample by simply combining a series of transmission wavefront error measurements using linearly polarized light with Fizeau interferometers. With this method, the birefringence inhomogeneity of two KAGRA’s input test masses with diameter of 22 cm is fully reconstructed. The birefringence information is then used to calculate the transverse beam shape of the light fields in orthogonal polarization directions when passing through the substrate. The calculated beam shapes show a good match with in-situ measurements using KAGRA interferometer. This technique is crucial for birefringence characterization of test masses in future detectors where even larger size are used.
@@ I Introduction
Thermal noise is one of the main noise sources for future gravitational wave detectors [1, 2]. To reduce thermal noise, future detectors like LIGO Voyager [3] and Einstein Telescope [4, 5] are designed to operate at cryogenic temperatures. Cosmic Explorer [6] also considers cryogenic operation for an upgrade. KAGRA [7, 8, 9, 10] is the only one operating at cryogenic temperatures and pioneering this technique among current large scale ground-based detectors. All these detectors apply large size high-quality mirrors as their test masses forming optical resonators to sense the tiny length fluctuations introduced by gravitational waves. Fused silica is commonly used to make test masses at room temperature, but it is not compatible with low temperatures. Crystalline materials are considered most promising to make cryogenic test masses. Sapphire was selected as the material for KAGRA’s test masses [11] due to its excellent thermal and optical properties at low temperature [12, 13]. However, recent studies reveal that birefringence of crystalline substrates and coatings can degrade the performance of the detector [14, 15, 16, 17, 18, 19]. In this case, the birefringence fluctuates in magnitude and direction throughout the substrate of the mirror. Therefore, it is necessary to measure the position dependence of the birefringence over a large area to assess its impact on the optical performance.
@@
There are many studies on two-dimensional birefringence measurements for transparent materials [14, 15, 20, 21, 22, 23, 24]. These measurements are either using a single-beam scanning technique, or expanding the beam to a large aperture and imaging with a sensor array. The scanning method is rather time-consuming for large size samples, e.g., mirrors with diameters usually more than 20 cm for ground-based gravitational wave detectors, if one wants to obtain good spatial resolution [25, 26]. The phase-shifting method in digital interferometry has been previously proposed for fast 2D birefringence measurements with a very high accuracy [27, 28, 29]. Fizeau interferometers are one of the most accurate methods to extract the phase information of an optical component (reflection phase or transmission phase) over a large area. Transmission Wavefront Error (TWE) is often used to measure the refractive index homogeneity of a substrate under test. In the presence of birefringence, however, a simple TWE measurement does not reveal the entire information about the substrate.
@@
In this paper we describe a method of combining a series of TWE map measurements using linearly polarized light in an appropriate way to fully reconstruct the refractive index information including birefringence. With this method, the birefringence inhomogeneity of two KAGRA’s input test masses (ITM) is extracted. This birefringence information is then used to calculate the transverse beam shape of the light fields in orthogonal polarization directions when passing through the substrate. The calculated beam shapes show a good match with in-situ beam shape measurements. The method of extracting birefringence information from TWE measurements with Fizeau interferometers is introduced in Section II. Birefringence characterization of KAGRA’s current ITMs is described in Section III. Our conclusions are summarized in Section IV.
@@ II Generating a birefringence map with a Fizeau interferometer
Figure 1 shows a schematic of a Fizeau interferometer. We analyze the interference pattern between the light reflected by the reference plane 1 ($E_{\mathrm{1}}$) and the light reflected by the reference plane 2 ($E_{\mathrm{2}}$) on the CCD. By comparing the results with and without the sample inserted into the cavity formed by the reference planes 1 and 2, we can obtain the information about the optical thickness of the sample. Since the optical thickness is the physical thickness times the index of refraction, by measuring the surface figure maps of S1 (surface with anti-reflecting coating) and S2 (surface with highly-reflective coating) precisely and subtracting out the variation of physical thickness of the sample, we can extract the fluctuation of the index of refraction in the substrate.
Figure 1: Schematic of TWE measurement with a Fizeau interferometer. S1 is the surface with anti-reflecting coating and S2 is the surface with highly-reflective coating. The interference between light fields $E_{\mathrm{1}}$ and $E_{\mathrm{2}}$ reflected by reference planes 1 and 2 respectively gives the TWE information of the sample.
In order to extract the birefringence information, measurements have to be performed using a linearly polarized light. To understand the behavior of such measurements, let us first consider what happens to a linearly polarized light going through a birefringent substrate.
II.1 Transmission of a linearly polarized light through a birefringent medium
Let us concentrate on an optical axis going through a single column of the substrate as shown in Figure 2. We define our preferred polarization axes called s and p in the vertical and horizontal directions with respect to the laboratory frame. The extra-ordinary and ordinary axes of the crystal in that particular column is assumed to be rotated by an angle $\theta$ with respect to our polarization axes. The input beam to the Fizeau interferometer is purely polarized in s ($E_{\mathrm{s}}$).
Figure 2: Transmission of a linearly polarized optical field through a column of a sample substrate.
We will use the Jones matrix formalism to analyze the conversion of the polarization state [30]. Here, the polarization state of light is represented by a vector,
@@ II.1 Transmission of a linearly polarized light through a birefringent medium
where $E_{\mathrm{s}}$ and $E_{\mathrm{p}}$ are the complex amplitudes of s- and p-polarized electric fields.
After passing through the substrate twice (a round trip through the substrate), the polarization state is converted to
@@
where $M$ is the round-trip Jones matrix of the substrate. We denote the indices of refraction for the extra-ordinary and ordinary axes as $n_{\mathrm{e}}$ and $n_{\mathrm{o}}$, respectively. Then we can compute the round trip phase changes of optical fields polarized in those axes:
@@
where $d$ is the thickness of the substrate, $\lambda$ is the wavelength of the light.
Using $\alpha_{\mathrm{e}}$ and $\alpha_{\mathrm{o}}$, we define the common and differential phase changes,
@@
For a purely s-polarized input field, $\vec{V}=(1,0)^{\mathrm{T}}$, the field returning from the substrate becomes,
@@
This $\vec{V}^{\prime}$ is the same as the field $E_{2}$ in Figure 1 at the CCD, except for the common phase term, which is ignored.
@@ II.2 Interference with the reference field
The field $\vec{V}^{\prime}$ computed above is superposed with the reference field ($E_{1}$ in Figure 1). The reference field $\vec{V}_{\mathrm{r}}$ is purely polarized in s and has a certain phase offset $\phi$ from $\vec{V}^{\prime}$,
@@
The optical power detected by the CCD after the interference of $\vec{V}^{\prime}$ and $\vec{V}_{\mathrm{r}}$ is,
@@
By plugging Equations 7 and 8 into Equation 9, we get,
@@
From the conservation of energy, $|E_{\mathrm{s}}|^{2}+|E_{\mathrm{p}}|^{2}=1$. From Equation 8, $|E_{\mathrm{r}}|^{2}=1$. We will ignore those constant terms in the following discussion. Then what is interesting to us is,
@@
If the s-polarization axis is aligned with the extra-ordinary axis ($\theta=0$), $\Delta P=2\cos(\phi-\alpha_{-})$. Usually, a Fizeau interferometer is equipped with a mechanism to change the position of the reference plane 1, for example, by mounting the reference flat on a piezo stage. Therefore, we can scan the relative phase $\phi$ by known amounts and fit the measured variation of $\Delta P$ to know the value of $\alpha_{-}$. In the general case of $\theta\neq 0$, the Fizeau measurement obtains the phase of the field $E_{\mathrm{s}}$, i.e. $\psi=\arg(E_{\mathrm{s}})$.
@@ II.3 Extraction of $\alpha_{-}$ and $\theta$ from several TWE maps
When the orientation of the extra-ordinary axis is not known, which is almost always the case, we need to take several TWE maps with different orientation of the sample or the measurement beam polarization. Then combine them to extract $\theta$ and $\alpha_{-}$.
As explained in the previous section, a TWE measurement by a Fizeau interferometer yields a transmission phase $\psi$ at each point on the sample surface. Let us assume that we rotate the sample by an angle $\eta$, or equivalently the input polarization orientation by $-\eta$. This is equivalent to changing the orientation of the extra-ordinary axis from $\theta$ to $\theta+\eta$.
Let us denote the measured phase with an $\eta$ rotation of the sample as $\psi(\eta)$. From Equation 7,
@@
We can see $\psi(\eta)$ as a function of $\eta$ with parameters $\theta$ and $\alpha_{-}$ determining its shape. Therefore, if we have several measurements of $\psi(\eta)$ with different values of $\eta$, we can use a non-linear fitting algorithm to fit the measured values with Equation 12 to obtain the best fit values of $\theta$ and $\alpha_{-}$.
If $\alpha_{-}$ can be assumed to be very small, which is the case for KAGRA mirrors, we can expand Equation 12 with $\alpha_{-}$, leaving only the $\mathcal{O}(\alpha_{-})$ terms, as
@@
We then take the ratio of $\Delta\psi_{1}$ and $\Delta\psi_{2}$,
@@
Once the value of $\theta$ is obtained, we can use Equation 12 to compute the value of $\alpha_{-}$ as,
@@
III Birefringence characterization of KAGRA sapphire ITMs
Sapphire is a uniaxial crystal with optical axis called c-axis. KAGRA’s sapphire test masses are fabricated so that the input surface is perpendicular to the c-axis. Light propagating in the test mass substrate will travel along the c-axis and there should be no birefringence in theory. However, local impurities and inhomogeneities in the crystal, and also mechanical and thermal stress changes (e.g., due to the suspension system supporting the mirror) can cause depolarization. These local defects will distort the polarization distribution of the light field in the transverse plane. From Equation 2 we can see the cross-coupling term $E_{\mathrm{p}}=-i\sin\alpha_{-}\sin 2\theta$ can be eliminated when the differential phase $\alpha_{-}$ equals 00. In this case, there will be no depolarization at your input polarization. However, gravitational wave detectors have very strict requirements on the TWE of light going through input test mass substrates which is determined by $\alpha_{+}$ in Equation 2. To achieve this requirement and to minimize the influence of crystal inhomogeneity, ion beam figuring technique is used during the polishing process in which the thickness of the substrate is precisely adjusted over the surface to make sure the TWE rms is within a few nanometers [18]. This results in the consequence that $\alpha_{+}$ and $\alpha_{-}$ of the wavefront distortions due to inhomogeneity and birefringence cannot be minimized simultaneously. The wavefront error and depolarization is a trade off in birefringent medium.
@@ III.1 TWE of KAGRA sapphire test mass
Figure 3: TWE maps of KAGRA’s current ITMY within the 140 mm aperture measured with linearly polarized light with different $\theta$.
The sapphire ITMs have been installed in KAGRA and directly birefringence mapping is impossible. Characterization of the current sapphire test masses has been studied in [31]. TWE maps of ITMs were measured several times with linearly polarized light by rotating the sample at different angles, which is equivalent with measurements by rotating the input polarization. TWE maps were supposed to be measured at ${}^{\circ}$, ${}^{\circ}$, ${}^{\circ}$ and ${}^{\circ}$. Methods described in Section II have a rigid requirement on the angular control. However, the measurements in [31] were not optimized for birefringence characterization. So, the rotation angle were not precisely controlled at that time. By comparing locations of the markers on the edge of the mirror in different maps, we were able to recover the real rotation angle (shown in Figure 3). One should notice that TWEs of KAGRA current ITMs measured with linearly polarized light are almost one-order of magnitude larger than specifications, because the vendor used circularly polarized light for ion beam figuring during polishing process. It can be seen from Figure 3, the ${}^{\circ}$ map and the ${}^{\circ}$ map have almost opposite values (also works for ${}^{\circ}$ and ${}^{\circ}$ maps). If summing two maps together, one can derive the TWE for circularly polarized light, which has very small residuals.

### v2

# Characterisation of birefringence inhomogeneity of KAGRA sapphire mirrors from transmitted wavefront error measurements

@@ Abstract
Cooling down test masses to cryogenic temperatures is a way to reduce the thermal noise of gravitational wave detectors. Crystalline materials are considered the most promising materials for fabricating cryogenic test masses and their coatings because of their excellent thermal and optical properties at low temperatures. However, birefringence owing to local impurities and inhomogeneities in the crystal can degrade the performance of the detector. The birefringence measurement or mapping over a two-dimensional area is thus important. This study describes a method for fast birefringence measurements of a large sample by simply combining a series of transmission wavefront error measurements using linearly polarised light with Fizeau interferometers. Using this method, the birefringence inhomogeneity information of KAGRA’s two input test masses with a diameter of 22 cm was fully reconstructed. The birefringence information was then used to calculate the transverse beam shapes of the light fields in orthogonal polarisation directions when passing through the substrate. The calculated beam shapes showed a good agreement with the in-situ measurements using the KAGRA interferometer. This technique is crucial for birefringence characterisation of test masses in future detectors, where even larger sizes are used.
@@ I Introduction
Thermal noise is one of the main noise sources for gravitational wave detectors [1, 2]. To reduce thermal noise, future detectors such as the LIGO Voyager [3] and Einstein Telescope [4, 5] are designed to operate at cryogenic temperatures. Cosmic Explorer [6] considers cryogenic operations for upgrades. KAGRA [7, 8, 9, 10] is the only detector operating at cryogenic temperatures and pioneers this technique among the current large-scale ground-based detectors. All these detectors use large, high-quality mirrors as their test masses, forming optical resonators to sense the tiny length fluctuations introduced by gravitational waves. Fused silica is commonly used to make test masses at room temperature; however, it is not compatible with low temperatures. Crystalline materials are considered to be the most promising materials for fabricating cryogenic test masses. Sapphire was selected as the material for KAGRA’s test masses [11] because of its excellent thermal and optical properties at low temperatures [12, 13]. However, recent studies have revealed that the birefringence of crystalline substrates and coatings can degrade detector performance [14, 15, 16, 17, 18, 19]. In this case, the birefringence fluctuates in magnitude and direction throughout the substrate of the mirror. Therefore, it is necessary to measure the position dependence of the birefringence over a large area to assess its impact on optical performance.
@@
There are several studies on two-dimensional birefringence measurements of transparent materials [14, 15, 20, 21, 22, 23, 24]. These measurements use either a single-beam scanning technique or expand the beam to a large aperture for imaging using a sensor array. The scanning method is rather time-consuming for large sample sizes, for example, mirrors with diameters usually greater than 20 cm for ground-based gravitational wave detectors, if one wants to obtain good spatial resolution [25, 26]. A phase-shifting method in digital interferometry has been proposed for fast 2D birefringence measurements with high accuracy [27, 28, 29]. Fizeau interferometry is among the most accurate methods for extracting the phase information of an optical component (reflection or transmission phase) over a large area. The Transmission Wavefront Error (TWE) is often used to measure the refractive index homogeneity of a substrate under test. However, in the presence of birefringence a simple TWE measurement does not reveal all the information about the substrate.
@@
In this study, we describe a method for combining a series of TWE map measurements using linearly polarised light in an appropriate way to fully reconstruct refractive index information, including birefringence. Using this method, the birefringence inhomogeneity information of KAGRA’s two input test masses (ITM) is extracted. This birefringence information was then used to calculate the transverse beam shape of the light fields in orthogonal polarisation directions when passing through the substrate. The calculated beam shapes were in good agreement with the in-situ beam shape measurements. The method of extracting birefringence information from TWE measurements with Fizeau interferometers is introduced in Section II. The birefringence characterisation of KAGRA’s current ITMs is described in Section III, and the conclusions are summarised in Section IV.
@@ II Generating a birefringence map with a Fizeau interferometer
Figure 1: Schematic of TWE measurement with a Fizeau interferometer. S1 is the surface with anti-reflection coating and S2 is the surface with highly reflective coating. The interference between light fields $E_{\mathrm{1}}$ and $E_{\mathrm{2}}$ reflected by reference planes 1 and 2, respectively, obtains the TWE information of the sample.
Figure 1 shows a schematic diagram of the Fizeau interferometer. We analyse the interference pattern between the light reflected by the reference plane 1 ($E_{\mathrm{1}}$) and the light reflected by the reference plane 2 ($E_{\mathrm{2}}$) on the CCD. By comparing the results obtained with and without the sample inserted into the cavity formed by reference planes 1 and 2, we obtained information on the optical thickness of the sample. Because the optical thickness is the physical thickness multiplied by the index of refraction, by measuring the surface figure maps of S1 (surface with anti-reflection coating) and S2 (surface with highly reflective coating) precisely and subtracting the variation in the physical thickness of the sample, we can extract the fluctuation of the index of refraction in the substrate.
To extract birefringence information, measurements must be performed using linearly polarised light. To understand the behaviour of such measurements, we first considered the behaviour of linearly polarised light passing through a birefringent substrate.
II.1 Transmission of a linearly polarised light through a birefringent medium
Consider optical axis passing through a single substrate column, as shown in Figure 2. We defined our preferred polarisation axes, s and p, in the vertical and horizontal directions with respect to the laboratory frame. The extraordinary and ordinary axes of the crystal in that particular column are assumed to be rotated by an angle $\theta$ with respect to our polarisation axes. The input beam to the Fizeau interferometer was purely polarised in s ($E_{\mathrm{s}}$).
Figure 2: Transmission of a linearly polarised optical field through a column of a sample substrate.
The Jones matrix formalism was used to analyse the conversion of the polarisation state [30]. The polarisation state of light is represented by the following vector:
@@ II.1 Transmission of a linearly polarised light through a birefringent medium
where $E_{\mathrm{s}}$ and $E_{\mathrm{p}}$ are the complex amplitudes of the s- and p-polarised electric fields, respectively. After passing through the substrate twice (a round trip through the substrate), the polarisation state was converted to:
@@
where $M$ denotes the round-trip Jones matrix of the substrate. The indices of refraction for the extraordinary and ordinary axes are denoted as $n_{\mathrm{e}}$ and $n_{\mathrm{o}}$, respectively. Then, we can compute the round- trip phase changes of the optical fields polarised along the axes:
@@
where $d$ is the thickness of the substrate, and $\lambda$ is the wavelength of the light.
Using $\alpha_{\mathrm{e}}$ and $\alpha_{\mathrm{o}}$, we define the common and differential phase changes as follows:
@@
For a purely s-polarised input field, $\vec{V}=(1,0)^{\mathrm{T}}$, the field returning from the substrate becomes,
@@
This $\vec{V}^{\prime}$ is the same as the field $E_{2}$, as shown in Figure 1, at the CCD, except for the common phase term, which is ignored.
@@ II.2 Interference with the reference field
The field $\vec{V}^{\prime}$ computed above is superposed with the reference field ($E_{1}$ in Figure 1). The reference field $\vec{V}_{\mathrm{r}}$ is purely polarised in s and has a certain phase offset $\phi$ from $\vec{V}^{\prime}$,
@@
The optical power detected by the CCD after the interference of $\vec{V}^{\prime}$ and $\vec{V}_{\mathrm{r}}$ is as follows:
@@
By substituting Equations (7) and (8) into Equation (9), we obtain:
@@
Based on energy conservation, $|E_{\mathrm{s}}|^{2}+|E_{\mathrm{p}}|^{2}=1$. From Equation (8), $|E_{\mathrm{r}}|^{2}=1$. We ignore these constant terms in the following discussion. However, note that
@@
If the s-polarisation axis is aligned with the extra-ordinary axis ($\theta=0$), $\Delta P=2\cos(\phi-\alpha_{-})$. A Fizeau interferometer is equipped with a mechanism to change the position of the reference plane 1, for example, by mounting a reference flat on a piezo stage. Therefore, we can scan the relative phase $\phi$ by known amounts and fit the measured variation of $\Delta P$ to obtain the value of $\alpha_{-}$. In the general case of $\theta\neq 0$, the Fizeau measurement yields the phase of the field $E_{\mathrm{s}}$, that is, $\psi=\arg(E_{\mathrm{s}})$.
@@ II.3 Extraction of $\alpha_{-}$ and $\theta$ from several TWE maps
When the orientation of the extraordinary axis is unknown, which is almost always the case, we must obtain several TWE maps with different orientations of the sample or measurement beam polarisation. Then combine them to extract $\theta$ and $\alpha_{-}$.
As explained in the previous section, a TWE measurement by a Fizeau interferometer yields a transmission phase $\psi$ at each point on the sample surface. Let us assume that we rotate the sample by an angle $\eta$, or equivalently the input polarisation orientation by $-\eta$. This is equivalent to changing the orientation of the extra-ordinary axis from $\theta$ to $\theta+\eta$.
Let us denote the measured phase with an $\eta$ rotation of the sample as $\psi(\eta)$. From Equation (7),
@@
We can see $\psi(\eta)$ as a function of $\eta$ with parameters $\theta$ and $\alpha_{-}$ determining its shape. Therefore, if we have several measurements of $\psi(\eta)$ with different values of $\eta$, we can use a non-linear fitting algorithm to fit the measured values with Equation (12) to obtain the best fit values of $\theta$ and $\alpha_{-}$.
If $\alpha_{-}$ can be assumed to be very small, which is the case for KAGRA mirrors, we can expand Equation (12) with $\alpha_{-}$, leaving only the $\mathcal{O}(\alpha_{-})$ terms, as follows:
@@
We then use the ratio of $\Delta\psi_{1}$ and $\Delta\psi_{2}$,
@@
Once the value of $\theta$ is obtained, we can use Equation (12) to compute the value of $\alpha_{-}$ as follows:
@@
III Birefringence characterisation of KAGRA sapphire ITMs
Table 1: Rms values of TWE specifications for the KAGRA test masses and measurements with a circular/linear polarisation beam within an aperture of 140 mm [31].
Circular Linear
Specification polarisation polarisation
ITMX $<$5 nm 4.09 nm 25.90 nm
ITMY $<$5 nm 4.61 nm 32.72 nm
Sapphire is a uniaxial crystal with an optical axis called the c-axis. KAGRA’s sapphire test masses are fabricated such that the input surface is perpendicular to the c-axis. Light propagating in the test mass substrate travels along the c-axis, and theoretically, should experience no birefringence. However, local impurities and inhomogeneities in the crystal, as well as mechanical and thermal stress (e.g. due to the suspension system supporting the mirror), can cause depolarisation. These local defects distorted the polarisation distribution of the light field in the transverse plane. From Equation (2), we can see that the cross-coupling term $E_{\mathrm{p}}=-i\sin\alpha_{-}\sin 2\theta$ can be eliminated when the differential phase $\alpha_{-}$ equals 00. In this case, there was no depolarisation for the input polarisation. Gravitational wave detectors have very strict requirements on the TWE of light going through input test mass substrates. To achieve this requirement and minimise the influence of crystal inhomogeneity, when the birefringence was ignored, an ion beam figuring technique was used during the polishing process, in which the thickness of the substrate was precisely adjusted over the surface to ensure that the TWE rms was within a few nanometres [18]. However, in a birefringent medium, $\alpha_{+}$ and $\alpha_{-}$ in Equation (2) are both functions of thickness. This results in the consequence that $\alpha_{+}$ and $\alpha_{-}$ of the wavefront distortions due to inhomogeneity and birefringence cannot be minimised simultaneously by ajusting the thickness of the substrate. Wavefront errors and depolarisation are trade-offs in a birefringent medium.
@@ III.1 TWE of KAGRA sapphire test mass
The sapphire ITMs have already been installed in KAGRA, and direct birefringence mapping is impossible now. However, characterisation of the sapphire test masses was studied in [31]. The TWE maps of the ITMs were measured several times with linearly polarised light by rotating the sample at different angles, which was equivalent to the measurements obtained by rotating the input polarisation. The TWE maps were measured at the angles of 0, 45, 90 and 135. The methods described in Section II require a rigid angular control. However, the main motivation of the measurements in [31] was to see the stress induced change in $\alpha_{+}$, not $\alpha_{-}$, by different mounting orientations. Therefore, the rotation angle was not precisely controlled. To mitigate the problem, we compared the locations of the markers on the edge of the mirrors in different maps. In this way, we were able to recover the real rotation angle (Figure 3). The TWEs of the KAGRA current ITMs measured with linearly polarised light are around 5 to 6 times larger than the specifications because the vendor used circularly polarised light to measure the path length difference and calculate the removal from surface S2 during the ion beam figuring process (rms values are listed in Table 1). As shown in Figure 3, the 0 map and the 88 map have almost opposite values (the same is true for 49 and 130 maps). By summing the two maps, we can derive the TWE for circularly polarised light, which has small residuals.
Figure 3: TWE maps of KAGRA’s current ITMY within the 140 mm aperture measured with linearly polarised light with different $\theta$.